OpenStudy (anonymous):

Find the axis of symmetry of the graph of the function. f(x) = 3x^2 + 12x + 7

OpenStudy (abb0t):

It's a parabola

OpenStudy (anonymous):

so how would I find the axis of symmetry?

OpenStudy (mathmate):

Try completing squares to transform f(x) into f(x)=a(x-h)^2+k then x=h is the line of symmetry (passing through the vertex).

OpenStudy (anonymous):

is the formyla -b/2a ?

OpenStudy (mathmate):

yes, you could use the formula as well, x=-b/2a

OpenStudy (anonymous):

So I got -2 when I used that formula.... then what do I do?

OpenStudy (mathmate):

"Try completing squares to transform f(x) into f(x)=a(x-h)^2+k then x=h is the line of symmetry (passing through the vertex)." "yes, you could use the formula as well, x=-b/2a"

OpenStudy (mathmate):

When attempting math problems, whenever in doubt, reread the question! Also recall that "an axis of symmetry" is a line.

OpenStudy (anonymous):

Oh haha wow thank you!

OpenStudy (mathmate):

yw!